Suppose you draw a card, put it back in the deck, and draw another one. What is the probability that the first card is a four and the second is a seven? (Enter your probability as a fraction.)

Answer by Boreal(15235)

There are 4 fours and 4 sevens
(4/52)(4/52)=(1/13)^2=1/169